# A halo model approach for mock catalogs of time-variable strong gravitational lenses

### v4

@@ 4.2 Mock catalog of lensed SNe
Next, we discuss the features of mock catalogs for lensed SNe with multiple images. As mentioned before, we set the effective survey period of LSST to be $t_{\mathrm{eff}}=2.5~{}\mathrm{yr}$ to generate the mock catalogs. The total number in the fiducial mock catalog is $236$ for lensed SNe, which is somewhat smaller than the result of $248$ from OM10+. Of these, $32$ lenses, or approximately $14\%$ of all lenses, are primarily produced by subhalos. On the other hand, the mock catalog with the Chabrier IMF includes $133$ lenses, and $15\%$ of these lenses are produced by subhalos. When we add a criterion of maximum image separation as $\Delta\theta\leq(2/3)\theta_{\mathrm{PSF}}$, the total number of observed lenses reduces to $197$ in the fiducial case, $100$ in the Chabrier IMF case, which should be compared with $178$ in the case of OM10+. Note that approximately one-third of the lensed SNe is of Type Ia, which is consistent with OM10+.
As in the case of lensed QSOs, our catalogs of lensed SNe basically consist of double (two images and one faint central image) and quad (four images and one faint central image) lenses. The fraction of quad lenses in the fiducial (Chabrier) model is approximately $23\%~{}(23\%)$, while the one in OM10+ is about $28\%$.
It should be noted that the expected number of lensed SNe (especially Type Ia) in future surveys have been investigated in previous studies other than OM10+ (e.g., Wojtak et al., 2019; Goldstein et al., 2019; Pierel et al., 2021; Sainz de Murieta et al., 2023; Dong et al., 2024). Basically, our results are consistent with these previous studies, but with some differences due to different selection criteria of lenses. For example, in Sainz de Murieta et al. (2023), the peak brightness of multiply lensed Type Ia SNe are counted as detectable lenses, while we adopt a selection criterion of 0.7 magnitude brighter than the limiting magnitude of the survey for the peak brightness of either the faintest image in a double lens system or the third brightest image in a quadruple lens system, as described in Sec. 4. Goldstein et al. (2019) performed pixel-level simulations of lensed SNe considering the LSST observing strategy and following selection criteria based on photometric classification by spectral template fitting.
@@ 4.3 Distributions of time delays for lensed QSOs and SNe
The distributions of time delays and magnitude differences for quad lenses derived from our fiducial mock catalogs are displayed in Fig. 15. Quad lenses exhibit additional complexity due to the presence of more independent image pairs compared with double lenses. However, in the three panels on the right of Fig. 15, we clearly see some general trends as follows. The first arrival image tends to be fainter than the second and third arrival images, i.e., $\Delta m_{21}<0$ and $\Delta m_{31}<0$, and also tends to be brighter than the fourth image, i.e., $\Delta m_{41}>0$. Hence for quad lenses, it is anticipated that the third brightest image is most likely to arrive first, which is in agreement with OM10 and supports our choice to make use of the third brightest image for the calculation of the magnification bias. This selection ensures the successful detection of the first three images to arrive. No significant differences are found in this distribution for quad lenses between lensed QSOs and SNe.
@@ 4.4 Distributions of highly magnified QSOs and SNe
In addition to mock catalogs of multiply imaged QSOs and SNe, we construct mock catalogs of highly magnified systems, which are defined by lens systems with their maximum magnifications larger than a threshold value. With this definition, highly magnified systems include both single and multiple image systems. Such mock catalogs will be useful for investigating magnification bias in searching for high-redshift objects (e.g., Keeton et al., 2005). Recently, Yue et al. (2023) reported that J0025–0145 is a lensing system that is not multiply imaged but has significant magnification, which is referred to as an intermediate lensing system. Investigating the statistics of highly magnified systems also helps in discovering such intermediate lensing systems.
We employ two thresholds for highly magnified systems, one with the maximum magnification larger than 3 and the other with the maximum magnification larger than 10. For lensed QSOs, we find that the number of highly magnified systems is about $5700$ ($\mu>3$) and $570$ ($\mu>10$) for the fiducial Salpeter IMF case and approximately $4900$ ($\mu>3$) and $470$ ($\mu>10$) for the Chabrier IMF case. For lensed SNe, the number of highly magnified systems is about $860$ ($\mu>3$) and $180$ ($\mu>10$) for the fiducial Salpeter IMF mock and approximately $640$ ($\mu>3$) and $130$ ($\mu>10$) for the Chabrier IMF.
We plot the fraction of the number of images for our fiducial lensed QSOs in Fig. 16. For $\mu>3$, the majority of systems are two- and single-image (i.e., intermediate lens) systems, while in $\mu>10$, the majority are two- and four-image systems. Yue et al. (2023) suggests that J0025–0145 has a high magnification of $\mu\gg 1$ despite several systematic uncertainties, including the errors in the supermassive black hole mass measurement and bolometric correction. According to our fiducial mock catalogs, the probability of finding an intermediate lensing system with such a high magnification, $\mu>10$, is not negligible, around $15\%$, and it is predicted that there are about 80 intermediate lensing systems in LSST. This result is broadly consistent with the earlier analysis by Keeton et al. (2005). We also find similar results for the lensed SNe, which are not shown here.
We next discuss properties of lens objects causing these highly magnified systems. Such discussions have recently been advanced primarily in the context of gravitationally lensed gravitational waves. There is still disagreement in the literature as to which lens populations are most important. For instance, Hilbert et al. (2008) and Smith et al. (2018) concluded that galaxy clusters are the most important lens population using ray tracing results from cosmological $N$-body simulations. Meanwhile, Robertson et al. (2020) showed that massive galaxies are the most important lens population by performing a similar test using cosmological hydrodynamical simulations.
@@ 4.5 Stellar mass fractions for lensed QSOs and SNe
where $\kappa_{\mathrm{lens,gal}}(\bm{\theta})$ and $\kappa_{\mathrm{tot}}(\bm{\theta})$ are the convergence from the primary lens galaxy, as defined in Eq. (33), and the total convergence at the image position of $\bm{\theta}$, respectively. While in this paper we do not consider the microlensing effect, the stellar mass fractions of individual images are quite important when adding microlensing to our mock catalog (Vernardos et al., 2024).
@@ 5 Comparing with currently known samples of lensed QSOs
In Sec. 4, we present mock catalogs of lensed QSOs and SNe for a 10-yr long LSST observation run. Here, we compare our mock catalogs with known lensed QSOs. For this purpose, we employ the statistical sample of lensed QSOs from the SDSS Quasar Lens Search (SQLS; Oguri et al., 2006), covering about the area of $8000~{}\mathrm{deg}^{2}$, provided by Inada et al. (2012), as well as the sample from Gaia (Gaia Collaboration, 2018), covering about the area of about $22000~{}\mathrm{deg}^{2}$.
@@ 5.1 Comparing with the SQLS lens sample
For a fair comparison with the SQLS statistical lensed QSO sample presented in Inada et al. (2012), we generate a subsample of our mock catalogs to include objects with image separations in the range of $1~{}\mathrm{arcsec}<\Delta\theta<20~{}\mathrm{arcsec}$, an $i$-band magnitude difference less than 1.25 mag for doubles, and an effective point-spread function (PSF) magnitudes of $i_{\mathrm{eff}}<19.1$ for QSOs with $2.20.6<z<$ 2.2. Here the effective PSF magnitude is calculated by the effective magnification $\mu_{\mathrm{eff}}$, which is defined as (Oguri et al., 2012)
@@
While the SQLS statistical sample contains 4 quad lenses out of a total of 26 lenses, our fiducial Salpeter IMF mock catalogs predict $32\pm 6$ lensed QSOs, including $6\pm 1$ quad lenses, where this interval corresponds to a $68\%$ confidence level from our mock catalogs of five LSST realizations. The numbers in our mock catalogs coincide with those in observations at the $2\sigma$ level. In Fig. 19, we show the distribution of the maximum image separation in our fiducial SQLS-like mock catalogs compared to the SQLS sample. It should be noted that the SQLS sample contains SDSS J1004+4112 (Inada et al., 2003; Oguri et al., 2004; Inada et al., 2005) with the maximum image separation of about $15$ arcsec; OM10+ cannot predict such a large-separation lens, as can be seen from Fig. 7. In contrast, our fiducial model indeed predicts such a large-separation lens in the SDSS. We also find that our fiducial model reproduces well the number of lensed QSOs for the whole image separation range.
In Fig. 19, we also plot the result from the Chabrier IMF mock catalogs for comparison. We find that the Chabrier IMF cannot reproduce the observation. Specifically, the Chabrier IMF mock catalogs predict a total of $9\pm 2$ lensed QSOs, of which $2\pm 1$ are expected to be quad lenses. This indicates that the abundance of lensed QSOs favors the Salpeter IMF over the Chabrier IMF. Interestingly, this result is in line with previous strong lensing studies favoring the Salpeter IMF rather than the Chabrier IMF (see e.g., Treu et al., 2010; Oguri et al., 2014; Sonnenfeld et al., 2015). We note that there are several strong lensing studies that prefer Chabrier IMF rather than Salpeter IMF (e.g. Smith and Lucey, 2013; Smith et al., 2015; Sonnenfeld et al., 2019), which implies the dependence of the stellar IMF on the stellar mass and redshift. Incorporating such mass and redshift dependence of the stellar IMF is left for future work.
Figure 19: The number distribution of lensed QSOs as a function of the maximum image separations between multiple images from subsamples of our mock catalogs to construct SQLS-like mock catalogs. The red line shows the result for our fiducial Salpeter IMF mock catalogs, while the blue line shows the result for the Chabrier IMF mock catalog. The grey histogram displays the number distribution of lensed QSOs in the SQLS statistical sample (Inada et al., 2012).
@@ 5.2 Comparing with the Gaia lens sample
Following the analysis in Lemon et al. (2019, 2023), we generate a subsample of our mock catalogs to include objects with image separations in the range of $1~{}\mathrm{arcsec}<\Delta\theta<4~{}\mathrm{arcsec}$, the QSO redshift of $z<4$, and $G$-band magnitudes of $G<20.7$ for both images for double lenses and for at least three bright images for quad lenses. In our analysis we remove lensed QSOs without any information on the magnitude of the second brightest image for double lenses and the third brightest image for quad lenses.
To convert $i$-band magnitudes of lensed QSO images to Gaia $G$-band magnitudes, we use the SDSS spectroscopic QSO catalog cross-matched to Gaia (Lyke et al., 2020) to determine a $G-i$ relation, ignoring the redshift dependence and the scatter in the $G-i$ relation. Specifically we adopt the offset of $G-i=0.15$.
The Gaia lens sample is constructed based on Gravitationally Lensed Quasar Database as well as the latest follow-up result presented in Lemon et al. (2023) for 53.9% of the full sky, which corresponds to the typical areas searched for lenses (Lemon et al., 2023). While the Gaia sample contains 215 lensed QSOs including 25 quad lenses, our fiducial Salpeter mock catalogs predict $205\pm 14$ lensed QSOs, including $29\pm 5$ quad lenses. This indicates that our mock catalogs are in excellent agreement with the Gaia lens sample. It should be emphasized that in this Gaia-like selection, the quad fraction increases from $\sim 10\%$ for the LSST lens sample to $\sim 15\%$. As discussed in Fig. 6, this difference of the quad fraction simply reflects the difference of the survey depth between Gaia and LSST.
In Fig. 20, we show the number distribution as a function of the maximum image separation for our subsample of mock catalogs and the Gaia lens sample. We find that the image separation distribution in the whole range of $1~{}\mathrm{arcsec}<\Delta\theta<4~{}\mathrm{arcsec}$ is reproduced well. In contrast, as in the case of SQLS, the Chabrier IMF mock catalogs predict a significantly smaller number of lensed QSOs in Gaia. Specifically, the Chabrier IMF mock catalogs predict a total of $102\pm 15$ lensed QSOs, of which $20\pm 4$ are expected to be quad lenses.
Figure 20: Similar to Fig. 19, but for lensed QSOs observed in Gaia compiled from Gravitationally Lensed Quasar Database and Lemon et al. (2023).
@@ 6 Conclusion
With this model, we have generated mock catalogs for lensed QSOs and SNe found in the baseline survey planned in LSST as described in Sec. 4. We found that the LSST will discover about 3500 lensed QSOs and 200 SNe with resolved multiple images assuming the Salpeter IMF as shown in Table 3. We have presented the distributions of lens, source, and image properties from the mock catalogs. In addition to mock catalogs of multiple-image lens systems, we also created mock catalogs of highly magnified systems, including both multiple-image and single-image lenses, and discussed their statistical properties. We summarize below the highlights of our results and discuss our results in comparison with previous work such as OM10+.
@@
  • We have found that the probability of detecting highly magnified single-image systems, which is sometimes referred to as an intermediate lensed QSO, with high magnification, $\mu>10$, like J0025–0145 discovered in Yue et al. (2023) is around $15\%$ among all the high-magnification gravitational lens events, which suggests that about 80 such systems for lensed QSOs will be present in LSST.
@@
We note that we have set a lower limit on the mass of halos to $M_{\mathrm{hh}}\gtrsim 10^{11}~{}M_{\odot}$ for main halos or $M_{\mathrm{sh}}\gtrsim 10^{10}~{}M_{\odot}$ for subhalos for numerical reasons. This means that the abundance of very small image separation lenses, such as sub-arcsecond lensed SNe discussed in Goobar et al. (2023), can be underestimated. We leave the detailed comparison of the abundance of very small-separation lenses with observations to future work.

### v5

@@ 4.2 Mock catalog of lensed SNe
Next, we discuss the features of mock catalogs for lensed SNe with multiple images. As mentioned before, we set the effective survey period of LSST to be $t_{\mathrm{eff}}=2.5~\mathrm{yr}$ to generate the mock catalogs. The total number in the fiducial mock catalog is $288$ for lensed SNe, which is somewhat larger than the result of $248$ from OM10+. Of these, $85$ lenses, or approximately $30\%$ of all lenses, are primarily produced by subhalos. On the other hand, the mock catalog with the Chabrier IMF includes $169$ lenses, and $32\%$ of these lenses are produced by subhalos. When we add a criterion of maximum image separation as $\Delta\theta\leq(2/3)\theta_{\mathrm{PSF}}$, the total number of observed lenses reduces to $234$ in the fiducial case, $127$ in the Chabrier IMF case, which should be compared with $178$ in the case of OM10+. Note that approximately one-third of the lensed SNe is of Type Ia, which is consistent with OM10+.
As in the case of lensed QSOs, our catalogs of lensed SNe basically consist of double (two images and one faint central image) and quad (four images and one faint central image) lenses. The fraction of quad lenses in the fiducial (Chabrier) model is approximately $24\%$ ($23\%$), while the one in OM10+ is about $28\%$.
It should be noted that the expected number of lensed SNe (especially Type Ia) in future surveys have been investigated in previous studies other than OM10+ (e.g., 2019MNRAS.487.3342W; 2019ApJS..243....6G; 2021ApJ...908..190P; 2023MNRAS.526.4296S; 2024arXiv240710470D). Basically, our results are consistent with these previous studies, but with some differences due to different selection criteria of lenses. For example, in 2023MNRAS.526.4296S, the peak brightness of multiply lensed Type Ia SNe are counted as detectable lenses, while we adopt a selection criterion of 0.7 magnitude brighter than the limiting magnitude of the survey for the peak brightness of either the faintest image in a double lens system or the third brightest image in a quadruple lens system, as described in Sec. 4. 2019ApJS..243....6G performed pixel-level simulations of lensed SNe considering the LSST observing strategy and following selection criteria based on photometric classification by spectral template fitting.
@@ 4.3 Distributions of time delays for lensed QSOs and SNe
The distributions of time delays and magnitude differences for quad lenses derived from our fiducial mock catalogs are displayed in Fig. 15. Quad lenses exhibit additional complexity due to the presence of more independent image pairs compared with double lenses. However, in the three panels on the right of Fig. 15, we clearly see some general trends as follows. The first arrival image tends to be fainter than the second and third arrival images, i.e., $\Delta m_{21}<0$ and $\Delta m_{31}<0$, and also tends to be brighter than the fourth image, i.e., $\Delta m_{41}>0$. Hence for quad lenses, it is anticipated that the third brightest image is most likely to arrive first, which is in agreement with 2010MNRAS.405.2579O and supports our choice to make use of the third brightest image for the calculation of the magnification bias. This selection ensures the successful detection of the first three images to arrive. No significant differences are found in this distribution for quad lenses between lensed QSOs and SNe.
@@ 4.4 Distributions of highly magnified QSOs and SNe
In addition to mock catalogs of multiply imaged QSOs and SNe, we construct mock catalogs of highly magnified systems, which are defined by lens systems with their maximum magnifications larger than a threshold value. With this definition, highly magnified systems include both single and multiple image systems. Such mock catalogs will be useful for investigating magnification bias in searching for high-redshift objects (e.g., 2005ApJ...621..559K). Recently, 2023AJ....165..191Y reported that J0025–0145 is a lensing system that is not multiply imaged but has significant magnification, which is referred to as an intermediate lensing system. Investigating the statistics of highly magnified systems also helps in discovering such intermediate lensing systems.
We employ two thresholds for highly magnified systems, one with the maximum magnification larger than 3 and the other with the maximum magnification larger than 10. For lensed QSOs, we find that the number of highly magnified systems is about $7459$ ($\mu>3$) and $764$ ($\mu>10$) for the fiducial Salpeter IMF case and approximately $6422$ ($\mu>3$) and $629$ ($\mu>10$) for the Chabrier IMF case. For lensed SNe, the number of highly magnified systems is about $1098$ ($\mu>3$) and $230$ ($\mu>10$) for the fiducial Salpeter IMF mock and approximately $822$ ($\mu>3$) and $170$ ($\mu>10$) for the Chabrier IMF.
We plot the fraction of the number of images for our fiducial lensed QSOs in Fig. 16. For $\mu>3$, the majority of systems are two- and single-image (i.e., intermediate lens) systems, while in $\mu>10$, the majority are two- and four-image systems. 2023AJ....165..191Y suggests that J0025–0145 has a high magnification of $\mu\gg 1$ despite several systematic uncertainties, including the errors in the supermassive black hole mass measurement and bolometric correction. According to our fiducial mock catalogs, the probability of finding an intermediate lensing system with such a high magnification, $\mu>10$, is not negligible, around $13\%$, and it is predicted that there are about 100 intermediate lensing systems in LSST. This result is broadly consistent with the earlier analysis by 2005ApJ...621..559K. We also find similar results for the lensed SNe, which are not shown here.
We next discuss properties of lens objects causing these highly magnified systems. Such discussions have recently been advanced primarily in the context of gravitationally lensed gravitational waves. There is still disagreement in the literature as to which lens populations are most important. For instance, 2008MNRAS.386.1845H and 2018MNRAS.475.3823S concluded that galaxy clusters are the most important lens population using ray tracing results from cosmological $N$-body simulations. Meanwhile, 2020MNRAS.495.3727R showed that massive galaxies are the most important lens population by performing a similar test using cosmological hydrodynamical simulations.
@@ 4.5 Stellar mass fractions for lensed QSOs and SNe
where $\kappa_{\mathrm{lens,gal}}(\bm{\theta})$ and $\kappa_{\mathrm{tot}}(\bm{\theta})$ are the convergence from the primary lens galaxy, as defined in Eq. (33), and the total convergence at the image position of $\bm{\theta}$, respectively. While in this paper we do not consider the microlensing effect, the stellar mass fractions of individual images are quite important when adding microlensing to our mock catalog (2024SSRv..220...14V).
@@ 5 Comparing with currently known samples of lensed QSOs
In Sec. 4, we present mock catalogs of lensed QSOs and SNe for a 10-yr long LSST observation run. Here, we compare our mock catalogs with known lensed QSOs. For this purpose, we employ the statistical sample of lensed QSOs from the SDSS Quasar Lens Search (SQLS; 2006AJ....132.. [^9] O), covering about the area of $8000~\mathrm{deg}^{2}$, provided by 2012AJ....143..119I, as well as the sample from Gaia (2018A&A...616A...1G), covering about the area of about $22000~\mathrm{deg}^{2}$.
@@ 5.1 Comparing with the SQLS lens sample
For a fair comparison with the SQLS statistical lensed QSO sample presented in 2012AJ....143..119I, we generate a subsample of our mock catalogs to include objects with image separations in the range of $1~\mathrm{arcsec}<\Delta\theta<20~\mathrm{arcsec}$, an $i$-band magnitude difference less than 1.25 mag for doubles, and an effective point-spread function (PSF) magnitudes of $i_{\mathrm{eff}}<19.1$ for QSOs with $0.6<z<2.2$. Here the effective PSF magnitude is calculated by the effective magnification $\mu_{\mathrm{eff}}$, which is defined as (2012AJ....143..120O)
@@
While the SQLS statistical sample contains 4 quad lenses out of a total of 26 lenses, our fiducial Salpeter IMF mock catalogs predict $43\pm 6$ lensed QSOs, including $9\pm 2$ quad lenses, where this interval corresponds to a $68\%$ confidence level from our mock catalogs of five LSST realizations. The numbers in our mock catalogs coincide with those in observations at the $3\sigma$ level. In Fig. 19, we show the distribution of the maximum image separation in our fiducial SQLS-like mock catalogs compared to the SQLS sample. It should be noted that the SQLS sample contains SDSS J1004+4112 (2003Natur.426..810I; 2004ApJ...605...78O; 2005PASJ...57L...7I) with the maximum image separation of about $15$ arcsec; OM10+ cannot predict such a large-separation lens, as can be seen from Fig. 7. In contrast, our fiducial model indeed predicts such a large-separation lens in the SDSS. We also find that our fiducial model reproduces well the number of lensed QSOs for the whole image separation range.
In Fig. 19, we also plot the result from the Chabrier IMF mock catalogs for comparison. We find that the Chabrier IMF cannot reproduce the observation. Specifically, the Chabrier IMF mock catalogs predict a total of $14\pm 2$ lensed QSOs, of which $4\pm 1$ are expected to be quad lenses. This indicates that the abundance of lensed QSOs favors the Salpeter IMF over the Chabrier IMF. Interestingly, this result is in line with previous strong lensing studies favoring the Salpeter IMF rather than the Chabrier IMF (see e.g., 2010ApJ...709.1195T; 2014MNRAS.439.2494O; 2015ApJ...800...94S). We note that there are several strong lensing studies that prefer Chabrier IMF rather than Salpeter IMF (e.g. 2013MNRAS.434.1964S; 2015MNRAS.449.3441S; 2019A&A...630A..71S), which implies the dependence of the stellar IMF on the stellar mass and redshift. Incorporating such mass and redshift dependence of the stellar IMF is left for future work.
Figure 19: The number distribution of lensed QSOs as a function of the maximum image separations between multiple images from subsamples of our mock catalogs to construct SQLS-like mock catalogs. The red line shows the result for our fiducial Salpeter IMF mock catalogs, while the blue line shows the result for the Chabrier IMF mock catalog. The grey histogram displays the number distribution of lensed QSOs in the SQLS statistical sample (2012AJ....143..119I).
@@ 5.2 Comparing with the Gaia lens sample
Following the analysis in 2019MNRAS.483.4242L; 2023MNRAS.520.3305L, we generate a subsample of our mock catalogs to include objects with image separations in the range of $1~\mathrm{arcsec}<\Delta\theta<4~\mathrm{arcsec}$, the QSO redshift of $z<4$, and $G$-band magnitudes of $G<20.7$ for both images for double lenses and for at least three bright images for quad lenses. In our analysis we remove lensed QSOs without any information on the magnitude of the second brightest image for double lenses and the third brightest image for quad lenses.
To convert $i$-band magnitudes of lensed QSO images to Gaia $G$-band magnitudes, we use the SDSS spectroscopic QSO catalog cross-matched to Gaia (2020ApJS..250....8L) to determine a GiG-i relation, ignoring the redshift dependence and the scatter in the GiG-i relation. Specifically we adopt the offset of Gi=0.15G-i=0.15.
The Gaia lens sample is constructed based on lensed_qso_database_clemon as well as the latest follow-up result presented in 2023MNRAS.520.3305L for 53.9% of the full sky, which corresponds to the typical areas searched for lenses (2023MNRAS.520.3305L). While the Gaia sample contains 215 lensed QSOs including 25 quad lenses, our fiducial Salpeter mock catalogs predict $306\pm 25$ lensed QSOs, including $51\pm 5$ quad lenses. This indicates that our mock catalogs are in agreement with the Gaia lens sample. It should be emphasized that in this Gaia-like selection, the quad fraction increases from $\sim 10\%$ for the LSST lens sample to $\sim 17\%$. As discussed in Fig. 6, this difference of the quad fraction simply reflects the difference of the survey depth between Gaia and LSST.
In Fig. 20, we show the number distribution as a function of the maximum image separation for our subsample of mock catalogs and the Gaia lens sample. We find that the image separation distribution in the whole range of $1~\mathrm{arcsec}<\Delta\theta<4~\mathrm{arcsec}$ is reproduced well. In contrast, as in the case of SQLS, the Chabrier IMF mock catalogs predict a significantly smaller number of lensed QSOs in Gaia. Specifically, the Chabrier IMF mock catalogs predict a total of $147\pm 6$ lensed QSOs, of which $27\pm 6$ are expected to be quad lenses.
Figure 20: Similar to Fig. 19, but for lensed QSOs observed in Gaia compiled from lensed_qso_database_clemon and 2023MNRAS.520.3305L.
@@ 6 Conclusion
With this model, we have generated mock catalogs for lensed QSOs and SNe found in the baseline survey planned in LSST as described in Sec. 4. We found that the LSST will discover about 4000 lensed QSOs and 200 SNe with resolved multiple images assuming the Salpeter IMF as shown in Table 3. We have presented the distributions of lens, source, and image properties from the mock catalogs. In addition to mock catalogs of multiple-image lens systems, we also created mock catalogs of highly magnified systems, including both multiple-image and single-image lenses, and discussed their statistical properties. We summarize below the highlights of our results and discuss our results in comparison with previous work such as OM10+.
@@
  • We have found that the probability of detecting highly magnified single-image systems, which is sometimes referred to as an intermediate lensed QSO, with high magnification, $\mu>10$, like J0025–0145 discovered in 2023AJ....165..191Y is around $13\%$ among all the high-magnification gravitational lens events, which suggests that about 100 such systems for lensed QSOs will be present in LSST.
@@
We note that we have set a lower limit on the mass of halos to $M_{\mathrm{hh}}\gtrsim 10^{10}~M_{\odot}$ for main halos or $M_{\mathrm{sh}}\gtrsim 10^{9}~M_{\odot}$ for subhalos for numerical reasons. This means that the abundance of very small image separation lenses, such as sub-arcsecond lensed SNe discussed in 2023NatAs...7.1098G, can be underestimated. We leave the detailed comparison of the abundance of very small-separation lenses with observations to future work.